OpenStudy (knightrider):

x^2 cos(xy) integrate?

OpenStudy (anonymous):

with respect to what x or y

OpenStudy (anonymous):

You can to integration by parts, however I prefer the Tabular Method: let u = x^2 v= cos(x) (until 0 is obtained) Alternating Signs U and its derivatives V and its antiderivatives + x^2 cos(x) - 2x sin(x) + 2 - cos(x) - 0 - sin (x) Now multiply diagonal terms and add: int. (x^2cos(x) = + x^2sin(x) + 2xcos(x) - 2sin(x) + C The Tabular Method can also be used for a few other integration problems. Try looking it up

OpenStudy (knightrider):

with respect to y

OpenStudy (anonymous):

nm i didnt see the xy*

OpenStudy (anonymous):

\[xsin(xy)\]

OpenStudy (knightrider):

how did u find it?

OpenStudy (anonymous):

treat x as a constant (for example pretend it is a 2) use then use u substitution

OpenStudy (anonymous):

do you understand

OpenStudy (knightrider):

yeah i got u thnx alot